Find the equivalent fraction of 3 / 5 having
numerator 9 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

We require numerator 9
Let N be the denominator of the fraction
∴9 / N =3 / 5
3 ×N =9 ×5
N = (9 × 5) / 3
= 15
Therefore the required fraction is 9 / 15

Thanks